### 10 Jan 2024 v1

# A proof of Audenaert-Kittaneh’s Conjecture

Abstract
We show that Audenaert-Kittaneh’s Conjecture related to $p$-Schatten class.
keywords:
Schatten $p$-norm, inequality
1 Introduction
The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathemetical physics. They have been generalised in various directions. Among these versions for more-general symmetric norms BH88 , for operators by $n$-tuples BK04 ; HK08 and for the Haagerup $L_{p}$ spaces FK86 , as well as refinements BCL94 . Let $\mathbb{B}(\mathscr{H})$ denote the algebra of all bounded linear operators acting on a complex separable Hilbert space $\mathscr{H}$. If $X\in\mathbb{B}(\mathscr{H})$ is compact, we denote by $\{s_{j}(X)\}$ the sequence of decreasingly ordered singular values of $X$. For $p>0$, let
$\|X\|_{p}=\left[(s_{j}(X))^{p}\right]^{1/p}=\left(\text{tr }\left|X\right|^{p}\right)^{\frac{1}{p}},$
[…]
$\mathbb{B}_{p}(\mathscr{H})=\{X\in\mathbb{B}(\mathscr{H}):\|X\|_{p}<\infty\},$
which is called the $p$-Schatten class of $\mathbb{B}(\mathscr{H})$; cf. GK69 .
Clarkson’s inequalities for operators in $\mathbb{B}_{p}(\mathscr{H})$ (see Mc67 ) assert that for $0<p\leq 2$
$2^{p-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p})\leq\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p}),$ (1)
and for $p\geq 2$
$2(\|A\|_{p}^{p}+\|B\|_{p}^{p})\leq\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2^{p-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p}).$ (2)
For $p=2$ both inequalities (1) and 2 reduce to the parallelogram law
$\|A-B\|_{2}^{2}+\|A+B\|_{2}^{2}=2(\|A\|_{2}^{2}+\|B\|_{2}^{2}).$
On the other hand, McCarthy Mc67 obtained for $1<p\leq 2$ the following inequality
$2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}\leq\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}$ (3)
and for $p\geq 2$
$\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}$ (4)
where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
One natural generalizations of (1) and (2) are as follows:
Theorem 1.1.
(HK08, , Hirazallah & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $0<p\leq 2$
$n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq\sum_{i=1}^{n}\|A_{i}\|^{p}_{p},$ (5)
and for $p\geq 2$
$n\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (6)
In view of the inequalities (3) and (4), it is reasonable to make the following conjecture concerning generalizations (along the lines of the inequalities (5) and (6) of the inequalities (1) and (2)) to $n$-tuples of operators.
Conjecture 1.2.
(AK12, , Audenaert & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\right)^{\frac{q}{p}}$
and for $p\geq 2$
$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
In CM16 , Conde and Moslehian obtained a weaker version of Conjecture 1.2 with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$, the conjecture 1.2 remains open.
Theorem 1.3.
(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then
$n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
for $p\in[2,\infty);\frac{1}{p}+\frac{1}{q}=1$,
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}$
for $p\in(1,2];\frac{1}{p}+\frac{1}{q}=1$.
In this paper, we give a proof of Audenaert-Kittaneh’s Conjecture along a completely different path from Conde and Moslehian.
2 Proof of Audenaert-Kittaneh’s Conjecture
In this section, we give a proof of Audenaert-Kittaneh’s Conjecture 1.2, the ideas are the same as those used in FK86 ; BK04 .
Lemma 2.4 (Hadamard three-lines theorem, see U08 ).
Let $f(z)$ be a bounded function of $z=x+iy$ defined on the strip
$\{x+iy:a\leq x\leq b\},$
holomorphic in the interior of the strip and continuous on the whole strip. If
$M(x):=\sup\limits_{y}\left|f(x+iy)\right|,$
then $\log M(x)$ is a convex function on $[a,b]$. In other words, if $x=ta+(1-t)b$ with $0\leq t\leq 1$, then
$M(x)\leq M(a)^{t}M(b)^{1-t}.$
Lemma 2.5.
Let $A_{1},\ldots,A_{n}\in B_{p}(\mathscr{H})$ for some $1<p\leq 2$. Denote $B=\sum\limits_{k=1}^{n}A_{k},B_{i,j}=A_{i}-A_{j},1\leq i<j\leq n$. Let $Y,Y_{i,j}$ be operators in the dual class $B_{q}(\mathscr{H})$. Then
|tr ($\left|\text{tr }\left(YB+\sum_{1\leq i<j\leq n}Y_{i,j}B_{i,j}\right)\right|\leq n^{\frac{1}{q}}\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)^{\frac{1}{p}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{p}}.$ (7)
Proof.
Let $A_{k}=\left|A_{k}\right|W_{k},Y=V\left|Y\right|$ and $Y_{i,j}=V_{i,j}\left|Y_{i,j}\right|$ be right, left and left polar decompositions of $A_{k},Y$ and $Y_{i,j}$, respectively. Here, $W_{j},V$ and $V_{i,j}$ are partial isometries. We have $11/2\leq 1/p<$ 1. For the complex variable $z=x+iy$ with $1/2\leq x\leq 1$, let
$A_{k}(z)$ $=$ $\left|A_{k}\right|^{pz}W_{k};$
$B(z)$ $=$ $\sum\limits_{k=1}^{n}A_{k}(z);$
$B_{i,j}(z)$ $=$ $A_{i}(z)-A_{j}(z);$
$Y(z)$ $=$ $\|Y\|_{q}^{pz-q(1-z)}V\left|Y\right|^{q(1-z)};$
$Y_{i,j}(z)$ $=$ $\|Y_{i,j}\|_{q}^{pz-q(1-z)}V_{i,j}\left|Y_{i,j}\right|^{q(1-z)}.$
Note that $A_{k}(1/p)=A_{k},Y(1/p)=Y$ and $Y_{i,j}(1/p)=Y_{i,j}$. Let
$f(z)=\text{tr }\left(Y(z)B(z)+\sum_{1\leq i<j<n}Y_{i,j}(z)B_{i,j}(z)\right).$
The left-hand side of (7) is $\left|f(1/p)\right|$. We can estimate this if we have bounds for $\left|f(z)\right|$ at $x=1/2$ and $x=1$. If $x=1$, we have
|tr $\left|\text{tr }Y(z)A_{k}(z)\right|$ $=$ $\|Y\|_{q}^{p}\left|\text{tr }V\left|Y\right|^{-iqy}\left|A_{k}\right|^{p(1+iy)}W_{k}\right|;$
|tr $\left|\text{tr }Y_{i,j}(z)A_{k}(z)\right|$ $=$ $\|Y_{i,j}\|_{q}^{p}\left|\text{tr }V_{i,j}\left|Y_{i,j}\right|^{-iqy}\left|A_{k}\right|^{p(1+iy)}W_{k}\right|.$
Using the information that for any operator $T$,
|tr $\left|\text{tr }T\right|\leq\|T\|_{1}\text{ and }\|XTY\|\leq\|X\|_{\infty}\|T\|\|X\|_{\infty}$
[…]
|tr $\left|\text{tr }Y(z)A_{k}(z)\right|$$\leq$ $\|Y\|_{q}^{p}\left\|A_{k}\right\|_{p}^{p};$
|tr $\left|\text{tr }Y_{i,j}(z)A_{k}(z)\right|$$\leq$ $\|Y_{i,j}\|_{q}^{p}\left\|A_{k}\right\|_{p}^{p}.$
Hence
$\left|f(z)\right|$ $=$ |tr ($\left|\text{tr }\left(Y(z)B(z)+\sum_{1\leq i<j\leq n}Y_{i,j}(z)B_{i,j}(z)\right)\right|$ (8)
$=$ |tr ($\left|\text{tr }\left(Y(z)\sum_{k=1}^{n}A_{k}(z)+\sum_{1\leq i<j\leq n}Y_{i,j}(z)\left(A_{i}(z)-A_{j}(z)\right)\right)\right|$
$=$ |tr $\left|\text{tr }\sum_{k=1}^{n}\left(Y(z)-\sum\limits_{i=1}^{k-1}Y_{i,k}(z)+\sum\limits_{i=k+1}^{n}Y_{k,i}(z)\right)A_{k}(z)\right|$
$\leq$ $\sum_{k=1}^{n}\left(\left|\text{tr }Y(z)A_{k}(z)\right|+\sum\limits_{i=1}^{k-1}\left|\text{tr }Y_{i,k}(z)A_{k}(z)\right|+\sum\limits_{i=k+1}^{n}\left|\text{tr }Y_{k,i}(z)A_{k}(z)\right|\right)$
$\leq$ $\sum_{k=1}^{n}\left(\|Y\|_{q}^{p}+\sum\limits_{i=1}^{k-1}\|Y_{i,k}\|_{q}^{p}+\sum\limits_{i=k+1}^{n}\|Y_{k,i}\|_{q}^{p}\right)\|A_{k}\|_{p}^{p}$
$\leq$ $\left(\left\|Y\right\|_{q}^{p}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{q}^{p}\right)\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)$
when $x=1$. When $x=1/2$, the operators $A_{k}(z)$ and $Y_{i,j}(z)$ are in $\mathbb{B}_{2}(\mathscr{H})$ and
$\left|f(z)\right|$$\leq$ |tr $\left|\text{tr }Y(z)B(z)\right|+\sum_{1\leq i<j\leq n}\left|\text{tr }Y_{i,j}(z)B_{i,j}(z)\right|$
$\leq$ $\|Y(z)\|_{2}\|B(z)\|_{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}\|B_{i,j}(z)\|_{2}$
$\leq$ $\left(\left\|Y(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}\left(\left\|B(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|B_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}$
$=$ $n^{\frac{1}{2}}\left(\left\|Y(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{2}^{2}\right)^{\frac{1}{2}}.$
The equality at the last step is a consequence of Theorem 1.3, specialised to the case $p=2$. Note that when $x=1/2$, we have $\|A_{k}(z)\|_{2}^{2}=\|A_{k}\|_{p}^{p}$, and $\|Y_{i,j}(z)\|_{2}^{2}=\|Y_{i,j}\|_{q}^{p}$. Hence
$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{p}^{p}\right)^{\frac{1}{2}},$ (9)
when $x=1/2$. Let $M_{1},M_{2}$ be the right sides of (8), (9), repectively. Then by Lemma 2.4, we have, for $11/2\leq 1/p<$ 1,
$\left|f(1/p)\right|\leq M_{1}^{2(1/p-1/2)}M_{2}^{2(1-1/p)}.$
This gives (7). ∎
Proof of Audenaert-Kittaneh’s Conjecture: Now, let $B=U\left|B\right|,B_{i,j}=U_{i,j}\left|B_{i,j}\right|$ be a polar decomposition, and let
$Y=\left\|B\right\|_{p}^{q-p}\left|B\right|^{p-1}U^{*},Y_{i,j}=\left\|B_{i,j}\right\|_{p}^{q-p}\left|B_{i,j}\right|^{p-1}U_{i,j}^{*}.$
It is easy to see that
$\text{tr }YB=\left\|B\right\|_{p}^{q}=\left\|Y\right\|_{q}^{p},\text{tr }Y_{i,j}B_{i,j}=\left\|B_{i,j}\right\|_{p}^{q}=\left\|Y_{i,j}\right\|_{q}^{p}.$
So from Lemma 2.5, we get
$\|B\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|B_{i,j}\|_{p}^{q}\leq n^{\frac{1}{q}}\left(\sum_{i=1}^{n}\left\|A_{i}\right\|_{p}^{p}\right)^{1/p}\left(\|B\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|B_{i,j}\|_{p}^{q}\right)^{1/p}.$
This is the same as saying that
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},1<p\leq 2.$
This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument.
Acknowledgments
I heartily thank my advisor Minghua Lin for his guidance in my study and care in my life.